OpenStudy (anonymous):

I DONT HAVE MUCH TIME TO GET THIS DONE, AND WOULD REALLY APPRECIATE SOME HELP Match the reasons to the corresponding statement to complete the proof. Write the letter of the reason in the box. Here are the statments: Statements Reasons 1) 5(x + 3) + 9 = 59 1) 2) 5x + 15 + 9 = 59 2) 3) 5x + 24 = 59 3) 4) 5x + 0 = 35 4) 5) 5x = 35 5) 6) x = 7 6) Please give the reason for each statment, the choices are below to choose from:) thanks so

OpenStudy (anonymous):

A. Reflexive Property B. Symmetric Property C. Transitive Property D. Addition Property of Equality E. Subtraction Property of Equality F. Multiplication Property of Equality G. Division Property of Equality H. Substitution Property I. Additive Identity J. Multiplicative Identity K. Commutative Property L. Associative Property M. Distributive Property N. Given

OpenStudy (anonymous):

I know the first one is give, so I just need five more, please guys help!:(

OpenStudy (anonymous):

given*

myininaya (myininaya):

What property is this a(b+c)=ab+ac?

OpenStudy (anonymous):

algebraic proof. wiel u please help me?

OpenStudy (anonymous):

you need help or answers?

OpenStudy (anonymous):

i am working on this also, and all i know is number 1

OpenStudy (anonymous):

well answers would be great considering my assignment is already late:)

OpenStudy (anonymous):

come on guys please....

OpenStudy (anonymous):

1) 5(x + 3) + 9 = 59 given

OpenStudy (anonymous):

5(x + 3) = 5x + 15 what will you call that?

OpenStudy (anonymous):

do you know number 3 and 5? thats all i NEED

OpenStudy (anonymous):

15 + 9 = 24 what operation did i use?

OpenStudy (anonymous):

im not quite sure.

OpenStudy (anonymous):

what does the + sign do

OpenStudy (anonymous):

I just need 3 and 5 guys!:(

OpenStudy (anonymous):

for 3 I got D and for 5 I got I. r these right?

OpenStudy (anonymous):

its right, but im sure u didnt do it :p

OpenStudy (anonymous):

lol i know ha